OpenStudy (anonymous):

Simplify (sin theta- cos theta)^2+(sin theta +cos theta)^2 A 1 B 2 C sin^2 theta D cos ^2 theta

OpenStudy (luigi0210):

FOIL

OpenStudy (anonymous):

Its been a while since ive done that, and was always bad at it, mind helping?

OpenStudy (luigi0210):

\(sin\theta-cos\theta)^2=(sin\theta-cos\theta)*(sin\theta-cos\theta)\)

OpenStudy (anonymous):

so the answer would just be 1?

OpenStudy (anonymous):

why?? just apply \((a\pm b)^2= a^2 \pm 2ab +b^2\) and then simplify, don't guess, work and you get the answer, it is not 1

OpenStudy (anonymous):

I was thinking it was one because the two sides would cross each other out, but give me a second and ill do it your way :D

OpenStudy (anonymous):

Alright i lied, i have no idea how to do that.